How do I do this problem?
Problem: There are two naturally occurring isotopes of Chlorine. 35Cl has a mass of 34.9689amu. 37Cl has a mass of 36.9659amu. Determine the abundance of each isotope.

35Cl= %
37Cl= %

The Atomic Mass of Chlorine is 35.453

Answer by ikleyn(52834)

.
Go to the link

https://answers.yahoo.com/question/index?qid=20180909234750AAwkHUq

and read it there.


            What follows, is copied and pasted from that Internet page.

            I copied and pasted it for you, but I am not responsible for it . . .


According to the periodic table, the average mass of one mole of chlorine is 35.453 grams. 

Let x be the fraction of the chlorine that is number 35. 1  -  x is the fraction that is number 37. 


34.9689 * x + 36.9659 * (1  -  x) = 34.9689 * x + 36.9659  -  36.9659 * x 

34.9689 * x + 36.9659  -  36.9659 * x = 35.453 

1.997 * x = 1.5359 

x = 1.5359 รท 1.997 


The decimal fraction is approximately 0.769. This means 76.9% is 35 Cl and 23.1% is 37 Cl. I hope this is helpful for you.